# Optimal position-building strategies in Competition

### v1

@@ 8.3 The two-trader equilibrium strategies
Suppose $a$ and $b$ in competition with $b$ $\lambda$-scaled. In this section we derive closed-form expressions for $a$ and $b$ that jointly satisfy Eqs. 19 to 20. Solving these yields:
$a(t)$ $=-\frac{\left(1-e^{-\frac{\kappa t}{3}}\right)\left(-e^{\kappa/3}\left(e^{\kappa/3}+e^{2\kappa/3}+1\right)(\lambda+1)+(\lambda-1)e^{\frac{\kappa t}{3}}+(\lambda-1)e^{\frac{2\kappa t}{3}}+(\lambda-1)e^{\kappa t}\right)}{2\left(e^{\kappa}-1\right)}$ (23)
$b(t)$ $=\frac{\left(1-e^{-\frac{\kappa t}{3}}\right)\left(e^{\kappa/3}\left(e^{\kappa/3}+e^{2\kappa/3}+1\right)(\lambda+1)+(\lambda-1)e^{\frac{\kappa t}{3}}+(\lambda-1)e^{\frac{2\kappa t}{3}}+(\lambda-1)e^{\kappa t}\right)}{2\left(e^{\kappa}-1\right)\lambda}$ (24)
@@ 9.1 The multi-trader symmetric equilibrium strategies
$L_{i}(t)=\dot{a}_{i}(t)\,\sum_{j}\dot{a}_{j}(t)+a_{i}(t)\cdot\kappa\cdot\sum_{j}a_{j}(t)$ (25)
In the above equation the temporary impact from traders $i=1,\dots,n+1$ is $\sum_{i=1}^{n}\dot{a}_{i}(t)$ while the permanent impact is $\kappa\,\sum_{i=1}^{n}a_{i}(t)$. Now using the Euler-Lagrange equation Eq. 6 we obtain that for each trader $i$ we obtain a second order differential equation for
$\ddot{a}_{i}=-\frac{1}{2}\left(\sum_{j\neq i}\ddot{a}_{j}+\kappa\,\sum_{j\neq i}\dot{a}\right),\quad\text{$a_{i}(0)=0,a_{i}(1)=1$ and $a_{i}=a_{j}$ for all $i$}$ (26)
We will call Eq. 26 the multi-trader symmetric equilibrium equation in light of the fact we demand all strategies are equal. Assuming that all $n+1$ traders will trade identical strategies in equilibrium and solving Eq. 26 simultaneously for all traders yields setting $\lambda=n$:
$a(t)=\frac{e^{\frac{\lambda\cdot\kappa}{\lambda+2}}\left(1-e^{-\frac{\lambda\cdot\kappa}{\lambda+2}\cdot t}\right)}{e^{\frac{\lambda\cdot\kappa}{\lambda+2}}-1}$ (27)
We call $a(t)$ in Eq. 27 the multi-trader symmetric equilibrium strategy and we plot it in Figure 16. Each plot shows the multi-trader equilibrium for three values of the number of the total number of traders, labeled $\lambda$ and show values for a total of 10, 50 and 250 traders. As we move down the rows $\kappa$ varies from a low- to high-$\kappa$ regime (see Section 4.4). For the low $\kappa$-regime we see the equilibrium strategies are almost risk-neutral, but as $\kappa$ grows the strategies become increasingly eager.
Figure 16: Multi-trader equilibrium Eq. 27 for various values of $\kappa$ and $\lambda$ throughout.
Next we examine what happens when the number of traders grows extremely large, roughly speaking, modeling the situation in which trader $A$ competes with many traders each trading a single unit. This is equivalent to letting $\lambda\to\infty$ in which case $\lambda\kappa/(\lambda+2)\to\kappa$ and Eq. 27 becomes, in the limit:
@@ 10.1 What is my strategy the best-response to?
$L(t;a,b^{*})=(\dot{a}+\lambda\dot{b}^{*})\,\dot{a}+\kappa\,(a+\lambda b^{*})\,\dot{a}$ (29)
@@
with boundary conditions $b(0)=0,b(1)=1$. As a sanity check we can take the solutions to the two-trader equilibrium equations Eqs. 19 to 20 and the solutions Eqs. 23 to 24 and see that if we use the solution for $a(t)$ in the system of equations and then solve for $b^{*}(t)$ in the inverse problem Eq. 30 we obtain the same $b^{(}t)=b(t)$ where $b(t)$ is taken from Eq. 23. In some sense this is no surprise as this is precisely the meaning of the equations, however Eq. 30 truly is answers the question if $a(t)$ is $A$’s strategy what is the $b^{*}(t)$ for which $a(t)$ is the best response? This has some useful properties which we explain now.
@@ 11.1 Overview
11.2 Analysis of $A$’s strategy selection
In this there is a single trader, $B$, who is trading some strategy $\lambda\cdot b(t)$. Suppose that $B$ believes that it is best to trade an equilibrium strategy, how does $B$ determine its strategy? The issues is that it depends on what $A$ believes is happening. Among all the possibilities, let’s suppose that $A$ correctly deduces that in addition to their trading one unit, there will be an additional $\lambda$ units of the stock traded. Further suppose that $A$ boils the situation down to two possibilities, either there is one adversary (the actual situation) or there are $\lambda$ adversaries, each trading one unit. Depending on which of the two possibilities $A$ believes they will trade a different equilibrium strategy and therefore $B$’s best-response will vary as well. To illustrate these two possibilities, we analyze what $B$’s best response is in each case.
@@ 13 Parameter mis-estimation
13.1 Costs arising from errors in $\kappa$
We start by giving a numerical exposition of how changes in $\kappa$ impact the total cost of trading in two-trader equilibrium strategies, specifically for the unit strategy $a(t)$.
@@ 13.1 Costs arising from errors in $\kappa$
Another view of how changes in $\kappa$ impact strategies is given in Figure in which we explicitly plot the two-trader equilibrium strategies $a(t),b(t)$ (see Eqs. 23 to 23) for various levels of $\kappa$. Each plot shows $a(t)$ for values of $\kappa$ shifted by 25% along with $b(t)$ in the right panel and the difference between the two $a(t)$ strategies in the left panel.
@@ 14 Two-trader equilibrium with risk aversion
Holding risk in this concept refers to a general aversion to holding the stock during the course of building the position. This may arise, for example, because while the predominant possibility is that the stock will rise during the position-building process and then "jump" when the catalyst occurs, there is a small chance that some bad news will occur and cause the price to fall. In this section augment two-trader equilibrium strategies (Section LABEL:sec:equilibrium-strategies) with an aversion to holding risk, as in Section 6.1 and which is a key feature in [AC97] and [AC01].
@@ 14.1 Augmented equilibrium equations
These in turn yield the two-trader equilibrium equations with risk analogous to Eqs. 19 to 20:
@@ 14.2 Exploration of two-trader equilibrium with risk-aversion
We briefly provide some numerical examples of the trading strategies satisfying Eqs. 51 to 52. We refer to the solutions as the two-trader equilibrium strategies with risk-aversion for the remainder [^8] We do not present the equations here but see Appendix 16 for that Mathematica code that produced Figure 24.. In Figure 24 we plot the solutions to these equations for various levels of risk aversion and constant levels of $\lambda$ and $\kappa$ (see the caption for details). The plots show the expected tension that arises between risk-aversion and the desire to minimize total cost of trading.
@@ 16 Appendix: Mathematica code for two-trader equilibrium with risk
We used Wolfram Mathematica [Inc] to solve Eqs. 51 to 52 but did not provide the explicit solutions. This is because they are fairly complex and their specific forms do not readily provide insight into the nature of the solutions. For completeness we provide the code used to generate Figure 24. The other plots in Section 14 were produced similarly.

### v2

@@ 8.3 The two-trader equilibrium strategies
Suppose $a$ and $b$ in competition with $b$ $\lambda$-scaled. In this section we derive closed-form expressions for $a$ and $b$ that jointly satisfy Eqs. 22 to 23. Solving these yields:
$a(t;{\kappa,\!\lambda\!})$ $=-\frac{\left(1-e^{-\frac{\kappa t}{3}}\right)\left(-e^{\kappa/3}\left(e^{\kappa/3}+e^{2\kappa/3}+1\right)(\lambda+1)+(\lambda-1)e^{\frac{\kappa t}{3}}+(\lambda-1)e^{\frac{2\kappa t}{3}}+(\lambda-1)e^{\kappa t}\right)}{2\left(e^{\kappa}-1\right)}$ (26)
$b(t;{\kappa,\!\lambda\!})$ $=\frac{\left(1-e^{-\frac{\kappa t}{3}}\right)\left(e^{\kappa/3}\left(e^{\kappa/3}+e^{2\kappa/3}+1\right)(\lambda+1)+(\lambda-1)e^{\frac{\kappa t}{3}}+(\lambda-1)e^{\frac{2\kappa t}{3}}+(\lambda-1)e^{\kappa t}\right)}{2\left(e^{\kappa}-1\right)\lambda}$ (27)
Before moving on to the next section we observe what happens when $\lambda=1$.
$\begin{split}a(t)&=\frac{(1-e^{-\frac{\kappa t}{3}})(e^{\kappa/3}+e^{2\kappa/3}+e^{\kappa})}{e^{\kappa}-1}\end{split}$ (28)
@@ 9.1 The multi-trader symmetric equilibrium strategies
$L_{i}(t)=\left(\sum_{j}\dot{a}_{j}(t)\right)\dot{a}_{i}(t)+\kappa\cdot\left(\sum_{j}a_{j}(t)\right)\dot{a}_{i}(t)$ (29)
In the above equation the temporary impact from traders $i=1,\dots,n+1$ is $\sum_{i=1}^{n}\dot{a}_{i}(t)$ while the permanent impact is $\kappa\,\sum_{i=1}^{n}a_{i}(t)$. Using the Euler-Lagrange equation Eq. 6 applied to trader $i$, that is taking derivatives with respect to $a_{i}$ and $\dot{a}_{i}$, we obtain:
$\ddot{a}_{i}=-\frac{1}{2}\left(\sum_{j\neq i}\ddot{a}_{j}+\kappa\,\sum_{j\neq i}\dot{a}_{i}\right),\quad\text{$a_{i}(0)=0,a_{i}(1)=1$ and $a_{i}=a_{j}$ for all $i$}$ (30)
We will call Eq. 30 the multi-trader symmetric equilibrium equation. To solve for the symmetric equilibrium we must solve 30 simultaneously for $i=1,\dots,n$, assuming that $a_{i}=a_{j}$ for all $i,j$. The result is an expression for $a(t)$ in terms of $n$. Then for $n+1$ traders we obtain:
$a(t)=\frac{1-e^{-\frac{n\cdot\kappa}{n+2}\cdot t}}{1-e^{\frac{n\cdot\kappa}{n+2}}}$ (31)
We call $a(t)$ in Eq. 31 the multi-trader symmetric equilibrium strategy for $n$ traders. And to clarify the quantities, this means that there are $n$ traders implicitly trading $1$ unit each for a total of $n+1$ units traded. Note that for $n=2$ traders this becomes:
$a(t)=\frac{e^{\kappa/3}\left(1-e^{-\frac{\kappa t}{3}}\right)}{e^{k/3}-1}$ (32)
And now we note that
$\frac{e^{\kappa/3}}{e^{k/3}-1}=\frac{e^{\kappa/3}+e^{2\kappa/3}+e^{\kappa}}{e^{\kappa}-1}$ (33)
so that Eq. 32 is equal to:
$a(t)=\frac{(1-e^{-\frac{\kappa t}{3}})(e^{\kappa/3}+e^{2\kappa/3}+e^{\kappa})}{e^{\kappa}-1}$ (34)
and this is the same as Eq. 28. This shows that the two-trader equilibrium strategy for when each trader trades one unit ($\lambda=1$) is the same as the $n$-trader equilibrium strategy specialized to the case of two-traders.
In Figure 16 we plot these. Each plot shows the multi-trader equilibrium for three values of the number of the total number of traders, labeled $\lambda$ and show values for a total of 10, 50 and 250 traders. As we move down the rows $\kappa$ varies from a low- to high-$\kappa$ regime (see Section 4.4). For the low $\kappa$-regime we see the equilibrium strategies are almost risk-neutral, but as $\kappa$ grows the strategies become increasingly eager.
Figure 16: Multi-trader equilibrium Eq. 31 for various values of $\kappa$ and $n$ throughout.
Next we examine what happens when the number of traders grows extremely large, roughly speaking, modeling the situation in which trader $A$ competes with many traders each trading a single unit. This is equivalent to letting $n\to\infty$ in which case $n\kappa/(n+2)\to\kappa$ and Eq. 31 becomes, in the limit:
@@
and this simplifies to:
$a_{i}(t)=\frac{1-e^{-\kappa t}}{1-e^{-\kappa}},\qquad i=1,\dots n$ (36)
@@ 10.1 What is my strategy the best-response to?
$L(t,a,b^{*})=(\dot{a}+\lambda\dot{b}^{*})\,\dot{a}+\kappa\,(a+\lambda b^{*})\,\dot{a}$ (37)
@@
with boundary conditions $b(0)=0,b(1)=1$. As a sanity check we can take the solutions to the two-trader equilibrium equations Eqs. 22 to 23 and the solutions Eqs. 26 to 27 and see that if we use the solution for $a(t)$ in the system of equations and then solve for $b^{*}(t)$ in the inverse problem Eq. 38 we obtain the same $b^{(}t)=b(t)$ where $b(t)$ is taken from Eq. 26. In some sense this is no surprise as this is precisely the meaning of the equations, however Eq. 38 truly is answers the question if $a(t)$ is $A$’s strategy what is the $b^{*}(t)$ for which $a(t)$ is the best response? This has some useful properties which we explain now.
@@ 11.1 Overview
11.2 Analysis of trader A’s strategy selection
In this, we begin with a trader, $B$, who is trading some strategy $\lambda\cdot b(t)$. Suppose that $B$ believes that it is best to trade an equilibrium strategy, how does $B$ determine its strategy? The issues is that it depends on what $A$ believes is happening. Among all the possibilities, let’s suppose that $A$ correctly deduces that in addition to their trading one unit, there will be an additional $\lambda$ units of the stock traded. Further suppose that $A$ boils the situation down to two possibilities, either there is one adversary (the actual situation) or there are $\lambda$ adversaries, each trading one unit. Depending on which of the two possibilities $A$ believes they will trade a different equilibrium strategy and therefore $B$’s best-response will vary as well. To illustrate these two possibilities, we analyze what $B$’s best response is in each case.
@@ 13 Parameter mis-estimation
13.1 Costs arising from errors in the market impact coefficient
We start by giving a numerical exposition of how changes in the market impact coefficient $\kappa$ effect the total cost of trading in two-trader equilibrium strategies, specifically for the unit strategy $a(t)$.
@@ 13.1 Costs arising from errors in the market impact coefficient
Another view of how changes in $\kappa$ impact strategies is given in Figure in which we explicitly plot the two-trader equilibrium strategies $a(t),b(t)$ (see Eqs. 26 to 26) for various levels of $\kappa$. Each plot shows $a(t)$ for values of $\kappa$ shifted by 25% along with $b(t)$ in the right panel and the difference between the two $a(t)$ strategies in the left panel.
@@ 14 Two-trader equilibrium with risk aversion
Holding risk in this context refers to a general aversion to holding the stock during the course of building the position. This may arise, for example, because while the predominant possibility is that the stock will rise during the position-building process and then "jump" when the catalyst occurs, there is a small chance that some bad news will occur and cause the price to fall. In this section we augment the two-trader equilibrium strategies discussed in Section 8 by adding a term the penalizes aversion to holding risk, as discussed in Section 6.1 and which is a key feature of optimal execution models as in [AC97] and [AC01].
@@ 14.1 Augmented equilibrium equations
These in turn yield the two-trader equilibrium equations with risk analogous to Eqs. 22 to 23:
@@ 14.2 Exploration of two-trader equilibrium with risk-aversion
We briefly provide some numerical examples of the trading strategies satisfying Eqs. 59 to 60. We refer to the solutions as the two-trader equilibrium strategies with risk-aversion for the remainder [^8] We do not present the equations here but see Appendix 16 for that Mathematica code that produced Figure 24.. In Figure 24 we plot the solutions to these equations for various levels of risk aversion and constant levels of $\lambda$ and $\kappa$ (see the caption for details). The plots show the expected tension that arises between risk-aversion and the desire to minimize total cost of trading.
@@ 16 Appendix: Mathematica code for two-trader equilibrium with risk
We used Wolfram Mathematica [Inc] to solve Eqs. 59 to 60 but did not provide the explicit solutions. This is because they are fairly complex and their specific forms do not readily provide insight into the nature of the solutions. For completeness we provide the code used to generate Figure 24. The other plots in Section 14 were produced similarly.